OpenStudy (anonymous):

Convert to degrees. -29pi/15

hero (hero):

pi = 180

hero (hero):

-29(180)/15

OpenStudy (primeralph):



OpenStudy (mathstudent55):

\(2 \pi = 360^o \) \(\pi = 180^o\) \( - \dfrac{29 \pi}{15} \times \dfrac{180^o}{\pi} \)

OpenStudy (anonymous):

thank you that helped alot